# The iterated Dirichlet process and applications to Bayesian inference

### v1

@@ 3.4 Mixtures and noisy observations
Proof.
Let $A\in\mathcal{S}^{n}$ and $B,C\in\mathcal{M}_{1}$. Using Proposition 3.3, we have
$P(\lambda\in B,\bm{\eta}_{n}\in A,\alpha\in C)$ $=E[1_{C}(\alpha)1_{B}(\lambda)P(\bm{\eta}_{n}\in A\mid\lambda,\alpha)]$
$=E[1_{C}(\alpha)1_{B}(\lambda)\lambda^{n}(A)]$
$=E[1_{C}(\alpha)E[1_{B}(\lambda)\lambda^{n}(A)\mid\alpha]]$
$=E\bigg{[}1_{C}(\alpha)\int_{B}\nu^{n}(A)\,\mathcal{D}(\alpha,d\nu)\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)\int_{M_{1}}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\nu^{n}(d{\bm{x}}_{n})\,\mathcal{D}(\alpha,d\nu)\bigg{]}.$
On the other hand,
$E\bigg{[}\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}1_{A}(\bm{\eta}_{n})\,$ $1_{C}(\alpha)\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)E\bigg{[}1_{A}(\bm{\eta}_{n})\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}\;\bigg{|}\;\lambda,\alpha\bigg{]}\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\lambda^{n}(d{\bm{x}}_{n})\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)E\bigg{[}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\lambda^{n}(d{\bm{x}}_{n})\;\bigg{|}\;\alpha\bigg{]}\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)\int_{M_{1}}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\nu^{n}(d{\bm{x}}_{n})\,\mathcal{D}(\alpha,d\nu)\bigg{]}.$
Hence,
[…]
which implies $P(\lambda\in B\mid\bm{\eta}_{n},\alpha)=\mathcal{D}(\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B)$. ∎
@@
In particular, if $\alpha$ is not random, as in the setting of Theorem 3.2, so that $Y$ and $\lambda$ are conditionally independent given $\bm{\eta}_{n}$, then
@@
Let $B\in\mathcal{M}_{1}$. Then
@@
3.5 Understanding the IDP parameters
Let $\varpi\sim\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$ be an IDP and $\xi=\{\xi_{mn}\}_{m,n\in\mathbb{N}}$ an array of samples from $\varpi$. By this we mean there exists a sequence of random measures $\mu=\{\mu_{m}\}_{m\in\mathbb{N}}$ such that ${\mu\mid\varpi}\sim\varpi^{\infty}$ and ${\xi_{m}\mid\mu_{m}}\sim\mu_{m}^{\infty}$. Note that each $\mu_{m}$ is a function of $\xi_{m}$. More specifically, the random measure $\mu_{m}$ is the almost sure weak limit of the empirical distributions $n^{-1}\sum_{j=1}^{n}\delta_{\xi_{mj}}$. As noted in Section 1, we call the $\mu_{m}$’s the row distributions, and we call $\varpi$ the row distribution generator.
Since ${\xi_{m}\mid\mu_{m}}\sim\mu_{m}^{\infty}$ and $\mu_{m}\sim\mathcal{D}(\varepsilon\varrho)$, we may apply the results of Section 3.3. For example, by (3.4), we have $P(\xi_{mn}\in A)=\varrho(A)$. In other words, the base measure $\varrho$ is the prior distribution on the samples $\xi_{mn}$.
Similarly, by (3.6), we have
$P(\xi_{m,n+1}\in A\mid X_{mn})=\frac{\varepsilon}{\varepsilon+n}\varrho(A)+\frac{1}{\varepsilon+n}\sum_{j=1}^{n}\delta_{\xi_{mj}}(A).$
As in Section 3.3, if the row concentration $\varepsilon$ is small, then our prior expectation on the empirical distributions of the rows $\xi_{m}$ is that they are concentrated on only a few points in $S$. Conversely, if $\varepsilon$ is large, then we expect those empirical distributions to be spread out over many points.
By Theorem 2.3 and (3.6) we have
$P(\xi_{m+1,n}\in A\mid\bm{\mu}_{m})$ $=E[P(\xi_{m+1,n}\in A\mid\bm{\mu}_{m+1})\mid\bm{\mu}_{m}]$
$=E[\mu_{m+1}(A)\mid\bm{\mu}_{m}]$
$=\int_{M_{1}}\nu(A)\,\mathcal{L}(\mu_{m+1}\mid\bm{\mu}_{m};d\nu)$
$=\frac{\kappa}{\kappa+m}\int_{M_{1}}\nu(A)\,\mathcal{D}(\varepsilon\varrho,d\nu)+\frac{1}{\kappa+m}\sum_{i=1}^{m}\mu_{i}(A).$
Hence, (3.3) gives
$P(\xi_{m+1,n}\in A\mid\bm{\mu}_{m})=\frac{\kappa}{\kappa+m}\varrho(A)+\frac{1}{\kappa+m}\sum_{i=1}^{m}\mu_{i}(A).$
As above, if the column concentration $\kappa$ is small, then we expect the empirical distribution of the $\mu_{m}$’s to be concentrated on only a few points in $M_{1}$. In other words, we expect the $\mu$’s to take on only a handful of different possible values. Conversely, if $\kappa$ is large, then we expect the $\mu_{m}$’s to take on many different values.
Remark 3.6.
The IDP bears some superficial resemblance to what is called a hierarchical Dirichlet process. A hierarchical Dirichlet process is a mixture of Dirichlet processes on $S$ with mixing distribution $\mathcal{L}(\alpha)$, where $\alpha/\alpha(S)$ is itself a Dirichlet process (see [15, 16]).
An IDP is a Dirichlet process on $M_{1}$ whose base measure is the law of a Dirichlet process. A hierarchical Dirichlet process, on the other hand, is as a Dirichlet process on $S$ whose base measure is a Dirichlet process itself. These are actually quite different. In fact, the two processes have different state spaces. A hierarchical Dirichlet process takes values in $M_{1}$, whereas an IDP takes values in $M_{1}(M_{1})$.
For example, if we take $\kappa^{\prime}$ to be random and $\rho^{\prime}\sim\mathcal{D}(\varepsilon\varrho)$, then we can define a hierarchical Dirichlet process $\lambda$ by the conditional distribution ${\lambda\mid\kappa^{\prime},\rho^{\prime}}\sim\mathcal{D}(\kappa^{\prime}\rho^{\prime})$. Note that $\rho^{\prime}$ is an $M_{1}$-valued random variable.
In contrast, for the IDP, we take $\kappa$ and $\rho$ to be nonrandom and $\rho=\mathcal{D}(\varepsilon\varrho)$. In this case, $\rho$ is a nonrandom element of $M_{1}(M_{1})$. We then define an IDP $\varpi$ by the unconditional distribution $\varpi\sim\mathcal{D}(\kappa\rho)$.
@@ 4 Sequential imputation
Recall that our goal is to compute (1.2), which we can do by computing (1.3). As we saw in Section 1, it will be unreasonable to calculate (1.3) exactly, at least when $m<M$, since in that case we would need to simultaneously condition on multiple rows of observations. If, however, we condition on only one row of observations, then we can make progress. (See Section 5.3.) Suppose, for instance, that we can compute (1.4) for $1\leq m\leq M$. If we can use these instead of (1.3) to compute (1.2), then we can avoid the issues that arose in the exact calculation of (1.3).
By Theorem 2.4, we must find a way to use (1.4) to compute $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. If we do this via simulation, then we must find a way to use (1.4) to simulate $\bm{\mu}_{M}$ according to the conditional distribution $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$.
One approach would be to simulate $\mu_{1}$ according to the distribution $\mathcal{L}(\mu_{1}\mid X_{1N})$, and then use that simulated value to simulate $\mu_{2}$ according to $\mathcal{L}(\mu_{2}\mid X_{2N},\mu_{1})$, and so on. However, if we do that, then we would not be simulating $\bm{\mu}_{M}$ according to its correct conditional distribution, since our simulation of $\mu_{m}$ would not take into account observations from higher-numbered rows.
@@
This is the method of sequential imputation, first introduced in [11]. It is an application of the more general method of importance sampling that originated in [9]. In this section, we show how to apply sequential imputation to the IDP. This was done in [12] in the special case $S=\{0,1\}$. The work in [12] was based on the work in [11], which evidently assumed that the underlying random variables have densities with respect to some product measure. In Proposition 4.5, we show that this assumption is not satisfied by the IDP. In particular, the work in [11] is not sufficient to justify its use in [12]. In Theorem 4.7, we provide a new proof that justifies this approach in general, including the case $S=\{0,1\}$ treated in [12] as well as the general case treated here.
@@ 4.1 Importance sampling
[…]
@@
[…]
@@
$E[w(Y,Z^{*,1})\mid Y]=\int_{S}w(Y,z)\,\mathfrak{m}^{*}(Y,dz)=h(Y).$ (4.4)
Hence, $Ew(Y,Z^{*,1})=Eh(Y)<\infty$. Now let $f:S\to\mathbb{R}$ be bounded and measurable. By Proposition 2.1, we have
[…]
@@ 4.2 Effective sample size
On the other hand, by (2.2), if $\{Z^{k}\}_{k=1}^{\infty}$ are copies of $Z$ that are i.i.d. given $Y$, then
@@
and call this the effective sample size.
@@
If we approximate $E[W\mid Y]$ by the sample mean, $\overline{W}=K^{-1}\sum_{k=1}^{K}W_{k}$, and $\operatorname{Var}(W\mid Y)$ by the population variance, $\widetilde{S}^{2}=(K^{-1}\sum_{k=1}^{K}W_{k}^{2})-\overline{W}^{2}$, then we have $K_{e}\approx K_{e}^{\prime}$, where
$K_{e}^{\prime}=\frac{K}{1+\widetilde{S}^{2}/\overline{W}^{2}}=\frac{\big{(}\sum_{k=1}^{K}W_{k}\big{)}^{2}}{\sum_{k=1}^{K}W_{k}^{2}}$
On the other hand, if we use the sample variance, $S^{2}=K(K-1)^{-1}\widetilde{S}^{2}$, then we have $K_{e}\approx K_{e}^{\prime\prime}$, where
$K_{e}^{\prime\prime}=\frac{K}{1+S^{2}/\overline{W}^{2}}=\frac{K(K-1)}{K-1+K\widetilde{S}^{2}/\overline{W}^{2}}=\frac{K(K-1)}{K-1+K(K/K_{e}^{\prime}-1)}=\left(\frac{K-1}{K-K_{e}^{\prime}/K}\right)K_{e}^{\prime}.$
@@ 4.3 Sequential imputation and the simulation measure
We think of $Y$ as observed values and $Z$ as unobserved. In this sense, $Z$ is regarded as “missing data.” We wish to determine $\mathcal{L}(Z\mid Y)$. Suppose, however, that we are only able to determine $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$ for $1\leq m\leq M$. (By convention, a variable with a 00 subscript is omitted. Hence, when $m=1$, we have $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})=\mathcal{L}(Z_{1}\mid Y_{1})$.) We describe here a method of using $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$ to approximate $\mathcal{L}(Z\mid Y)$. This method is called sequential imputation and first appeared in [11].
@@
In Theorem 4.1, we have $Z^{*,k}\mid Y\sim\mathfrak{m}^{*}(Y)$. Hence,
$Z_{m}^{*,k}\mid Y,Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k}\sim\gamma_{m}^{*}(Y,Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k})=\gamma_{m}({\bm{Y}}_{m},Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k}).$ (4.5)
In other words, the simulated vector $Z^{*,k}=(Z_{1}^{*,k},\ldots,Z_{M}^{*,k})$ can be constructed sequentially using $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$, where in each step the missing data ${\bm{Z}}_{m-1}$ is imputed with the previously simulated values $Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k}$. To prove that Theorem 4.1 applies in this situation, we must find a weight function $w$ satisfying (4.1).
@@ 4.4 A simulation density
In other words, such functions are the marginal densities. By changing $f$ on a set of measure zero, we may assume $f\in[0,\infty)$ everywhere and if the value of a marginal density at a point is 00, then $f$ at that point is 00 for all values of the omitted arguments.
We use |||| to denote conditional densities. For example,
@@
As usual, we adopt the convention that a variable with a 00 subscript is omitted. For instance, if $m=1$, then $f({\bm{y}}_{m},{\bm{z}}_{m-1})=f(y_{1})$ and $f(z_{m}\mid{\bm{y}}_{m},{\bm{z}}_{m-1})=f(z_{1}\mid y_{1})$.
@@ 4.5 A proof using densities
[…]
@@
[…]
@@
4.6 The IDP has no density
We wish to apply sequential imputation to determine $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$, using the computable distributions (1.4). In this case, we would naturally take $Z_{m}=\mu_{m}$ and $Y_{m}=X_{mN}=(\xi_{m1},\xi_{m2},\ldots,\xi_{mN})$. In [12], the author did exactly this in the special case $S=\{0,1\}$.
Unfortunately, as we see below in Proposition 4.5, sequential imputation—as it is presented in Theorem 4.4—does not apply in this case. Namely, Assumption 4.3 is not satisfied. The vector $Z=\bm{\mu}_{m}$ has no joint density with respect to any product measure. Hence, the proof of Theorem 4.4, which is a rigorous presentation of the proof in [11], does not justify the use of sequential imputation in this setting. This includes not only the general setting that we are working with, but also the special case $S=\{0,1\}$ that was treated in [12].
Proposition 4.5.
If $\varrho$ is not a point mass, then there do not exist $\sigma$-finite measures $\mathfrak{n}_{1}$ and $\mathfrak{n}_{2}$ on $M_{1}$ such that $\mathcal{L}(\bm{\mu}_{2})\ll\mathfrak{n}_{1}\times\mathfrak{n}_{2}$.
Proof.
Suppose $\varrho$ is not a point mass. Assume that $\mathfrak{n}_{1}$ and $\mathfrak{n}_{2}$ are $\sigma$-finite measures on $M_{1}$ and $\mathcal{L}(\bm{\mu}_{2})\ll\mathfrak{n}_{1}\times\mathfrak{n}_{2}$. Let $A=\{\gamma\in M_{1}:\mathfrak{n}_{1}(\{\gamma\})>0\}$, so that $A$ is countable. Let $D=\{(\gamma,\gamma)\in M_{1}^{2}:\gamma\notin A\}$. Then
$(\mathfrak{n}_{1}\times\mathfrak{n}_{2})(D)=\int_{A^{c}}\mathfrak{n}_{1}(\{\gamma\})\,\mathfrak{n}_{2}(d\gamma)=0.$
Therefore, by hypothesis, $P(\bm{\mu}_{2}\in D)=0$. Since $\varrho$ is not a point mass, the measure $\rho=\mathcal{D}(\varepsilon\varrho)$ has no discrete component. Therefore, $P(\mu_{1}\in A)=\rho(A)=0$, since $A$ is countable. Hence, by [7, Theorem 5.4] and (3.6), we have
$P(\bm{\mu}_{2}\in D)=P(\mu_{1}\notin A,\mu_{1}=\mu_{2})=P(\mu_{1}=\mu_{2})=E[P(\mu_{1}=\mu_{2}\mid\mu_{1})]\\ =E[\mathcal{L}(\mu_{2}\mid\mu_{1};\{\mu_{1}\})]=E\bigg{[}\frac{\kappa}{\kappa+1}\rho(\{\mu_{1}\})+\frac{1}{\kappa+1}\delta_{\mu_{1}}(\{\mu_{1}\})\bigg{]}=\frac{1}{\kappa+1}>0,$
a contradiction. ∎

### v2

@@ 2.3 Mixtures of Dirichlet processes
In particular, if $\alpha$ is not random, as in (2.2), so that $Y$ and $\lambda$ are conditionally independent given $\bm{\eta}_{n}$, then
@@
Let $B\subseteq M_{1}$ be Borel measurable. Then
@@ 3 Sequential imputation
As discussed in Section 1, we aim to find a way to use (1.4) to compute $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. If we do this via simulation, then we must find a way to use (1.4) to simulate $\bm{\mu}_{M}$ according to the conditional distribution $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. One approach would be to simulate $\mu_{1}$ according to the distribution $\mathcal{L}(\mu_{1}\mid X_{1N})$, and then use that simulated value to simulate $\mu_{2}$ according to $\mathcal{L}(\mu_{2}\mid X_{2N},\mu_{1})$, and so on. However, if we do that, then we would not be simulating $\bm{\mu}_{M}$ according to its correct conditional distribution, since our simulation of $\mu_{m}$ would not take into account observations from higher-numbered rows.
@@
This is the method of sequential imputation, first introduced in [10]. It is an application of the more general method of importance sampling that originated in [8]. In Sections 3.1 and 3.2, we give a generalized formulation of importance sampling, along with a discussion of effective sample size. In Section 3.3, we lay out the definitions and constructions that are needed in sequential imputation. In Sections 3.4 and 3.5, we prove that sequential imputation leads asymptotically to the desired conditional expectation. The proof in Section 3.5 is a rigorous presentation of the proof in [10].
Unfortunately, as we will see in Section 4, this version of sequential imputation does not apply to the NDP. In Section 3.6, therefore, we provide a new proof under more general assumptions. This new result, given in Theorem 3.6, will allow us in Section 4 to apply sequential imputation to the NDP.
@@ 3.1 Importance sampling
$w(Y,\cdot)=h(Y)\,\frac{d\mathcal{L}(Z\mid Y)}{d\mathfrak{m}^{*}(Y)}\quad\text{a.s.}$ (3.1)
@@
$\mathcal{L}(\widetilde{Z}^{K}\mid Z^{*,1},\ldots,Z^{*,K},Y)\to\mathcal{L}(Z\mid Y)\quad\text{a.s.}$ (3.3)
@@
$\begin{split}E[w(Y,Z^{*})\mid Y]&=\int_{S}w(Y,z)\,\mathfrak{m}^{*}(Y,dz)\\ &=\int_{S}h(Y)\frac{d\mathcal{L}(Z\mid Y)}{d\mathfrak{m}^{*}(Y)}\mathfrak{m}^{*}(Y,dz)\\ &=h(Y)\int_{S}d\mathcal{L}(Z\mid Y;dz)\\ &=h(Y).\end{split}$ (3.4)
Hence, $Ew(Y,Z^{*})=Eh(Y)<\infty$. Now let $f:S\to\mathbb{R}$ be bounded and measurable. By the conditional law of large numbers, we have
$\frac{1}{K}\sum_{k=1}^{K}{w(Y,Z^{*})f(Z^{*,k})}\to E[w(Y,Z^{*})f(Z^{*})\mid Y]\quad\text{a.s.}$
@@ 3.2 Effective sample size
On the other hand, by the conditional law of large numbers, if $\{Z^{k}\}_{k=1}^{\infty}$ are copies of $Z$ that are i.i.d. given $Y$, then
@@
which is called the effective sample size.
@@
If we approximate $E[W\mid Y]$ by the sample mean, $\overline{W}=K^{-1}\sum_{k=1}^{K}W_{k}$, and $\operatorname{Var}(W\mid Y)$ by the population variance, $\widetilde{\sigma}^{2}=(K^{-1}\sum_{k=1}^{K}W_{k}^{2})-\overline{W}^{2}$, then we have $K_{e}\approx K_{e}^{\prime}$, where
$K_{e}^{\prime}=\frac{K}{1+\widetilde{\sigma}^{2}/\overline{W}^{2}}=\frac{\big(\sum_{k=1}^{K}W_{k}\big)^{2}}{\sum_{k=1}^{K}W_{k}^{2}}.$
On the other hand, if we use the sample variance, $\sigma^{2}=K(K-1)^{-1}\widetilde{\sigma}^{2}$, then we have $K_{e}\approx K_{e}^{\prime\prime}$, where
$K_{e}^{\prime\prime}=\frac{K}{1+\sigma^{2}/\overline{W}^{2}}=\frac{K(K-1)}{K-1+K\widetilde{\sigma}^{2}/\overline{W}^{2}}=\frac{K(K-1)}{K-1+K(K/K_{e}^{\prime}-1)}=\left(\frac{K-1}{K-K_{e}^{\prime}/K}\right)K_{e}^{\prime}.$
@@ 3.3 Sequential imputation and the simulation measure
We think of $Y$ as observed values and $Z$ as unobserved. In this sense, $Z$ is regarded as “missing data.” We wish to determine $\mathcal{L}(Z\mid Y)$. Suppose, however, that we are only able to determine $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$ for $1\leq m\leq M$. (By convention, a variable with a 0 subscript is omitted. Hence, when $m=1$, we have $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})=\mathcal{L}(Z_{1}\mid Y_{1})$.) We describe here a method of using $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$ to approximate $\mathcal{L}(Z\mid Y)$. This method is called sequential imputation and first appeared in [10].
@@
In Theorem 3.1, we have $Z^{*}\mid Y\sim\mathfrak{m}^{*}(Y)$. Hence,
$Z_{m}^{*}\mid Y,Z_{1}^{*},\ldots,Z_{m-1}^{*}\sim\gamma_{m}^{*}(Y,Z_{1}^{*},\ldots,Z_{m-1}^{*})=\gamma_{m}({\bm{Y}}_{m},Z_{1}^{*},\ldots,Z_{m-1}^{*}).$ (3.5)
In other words, the simulated vector $Z^{*}=(Z_{1}^{*},\ldots,Z_{M}^{*})$ can be constructed sequentially using $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$, where in each step the missing data ${\bm{Z}}_{m-1}$ is imputed with the previously simulated values $Z_{1}^{*},\ldots,Z_{m-1}^{*}$. To prove that Theorem 3.1 applies in this situation, we must find a weight function $w$ satisfying (3.1).
@@ 3.4 A simulation density
In other words, such functions are the marginal densities. By changing $f$ on a set of measure zero, we may assume $f\in[0,\infty)$ everywhere and if the value of a marginal density at a point is 0, then $f$ at that point is 0 for all values of the omitted arguments.
We use $|$ to denote conditional densities. For example,
@@
As usual, we adopt the convention that a variable with a 0 subscript is omitted. For instance, if $m=1$, then $f({\bm{y}}_{m},{\bm{z}}_{m-1})=f(y_{1})$ and $f(z_{m}\mid{\bm{y}}_{m},{\bm{z}}_{m-1})=f(z_{1}\mid y_{1})$.
@@ 3.5 A proof using densities
$\mathcal{L}(\widetilde{Z}^{K}\mid Z^{*,1},\ldots,Z^{*,K},Y)\to\mathcal{L}(Z\mid Y)\quad\text{a.s.}$ (3.6)
@@
$w(Y,\cdot)=f(Y)\,\frac{d\mathcal{L}(Z\mid Y)}{d\mathfrak{m}^{*}(Y)}\quad\text{a.s.}$